1. You just retired with $1m! You are now planning your retirement distributions over the next 30 years. You expect to earn 6% annually and you expect inflation to be 3% annually. To ensure your purchasing power is maintained, youve decided youd like to withdraw the same amount each year in terms purchasing power. Please provide me with the following information: 1) the real rate of interest (do not approximate), 2) the real and nominal amount of your withdrawals in years 1, 10, an of d 30. (8 Points)
Answer #1
Nominal Rate of return 6%
Inflation 3%
Real rate= ((1+6%)/(1+3%)-1)
Real rate= 2.9126%
To support inflation backing, annuity should be increased each year by 3%
PV of annuity for growing annuity
P = (PMT/(r-g)) x (1-((1+g)/(1 + r)) ^n)
Where:
P = the present value of an annuity stream
PMT = the dollar amount of each annuity payment
r = the effective interest rate (also known as the discount rate)
i=nominal Interest rate
n = the number of periods in which payments will be made
g= Growth rate
PV 1000000
R= 6%
Years 30
Growth rate 3%
1000000= (PMT/(6%-3%)) x (1-((1+3%)/(1 + 6%)) ^30)
1000000= (PMT/(6%-3%)) * 0.577389
(PMT/(6%-3%))= 1000000/0.5777389
(PMT/(6%-3%))= 1,731,934.58
PMT=        51,958.04
So First annual payment by 51,958.04 which should increase by 3% every year
Nominal withdrawal Nominal withdrawal
T1        51,958.04     51,958.04
T10 51958.04*(1+3%)^9     67,793.46
T30 51958.04*(1+3%)^29 122,442.52
Real withdrawal Real withdrawal
T1 51958.04/(1+3%)^1     50,444.70
T10 67793.46/(1+3%)^10     50,444.70
T30 122442.52/(1+3%)^30     50,444.70